A surveyor measures the distance across a straight river by the following method:starting directly across from a tree on the opposite bank, he walks x=102m along the riverbank to establish a baseline. Then he sights across to the tree. The angle form his baseline to the tree is theta=39.8 degrees. How wide is the river ?


1
Expert's answer
2021-02-03T16:12:40-0500

Let "A" be the point of tree's location (see the figure). Then we need to find side "AB" from the right triangle "\triangle ABC". By definition of tan function, obtain:


"AB = x\tan \theta"

Substituting "x = 102 m, \theta = 39.8\degree", obtain:


"AB = 102\cdot \tan39.8\degree \approx 85m"

Answer. 85 m.